factorise x 3 -x-1

factorise x3 -x-1

Grade:10

5 Answers

Adarsh
8 years ago
x3−x-1=x(x2−x-2)=x(x-(1/x))(x+(1/x))
If it is correct please approve.
Adarsh
8 years ago
The formula used for solving this problem is
a2−b2=(a-b)(a+b)
ng29
8 years ago
@adarsh i think u had misunderrstood the question
the question is x3-x-1 and u have considered only x3-x ,
check it out
Adarsh
8 years ago
No my answer is right. I am solved it in the right way.
I u have any other answer please post it.
 
If my answer is correct please approve.
Gman Namg
8 years ago
first take x common and then take commons thus the answer would be like this
(x)(x+1)(x-1)-1
 
approve if useful